OpenStudy (anonymous):

how do you make (-1/2)-y=(-11/2)to standard form? show work please

OpenStudy (anonymous):

i only know there can be no fractions in standard form

OpenStudy (anonymous):

what if you just multiplied the equation by 2?

OpenStudy (anonymous):

i don't know how to do that

OpenStudy (anonymous):

ok so take the left hand side which is \[-\frac{1}{2}-y \] and multiply each term by 2

OpenStudy (anonymous):

so do 2times 1 then 2times 11?

OpenStudy (anonymous):

\[2\times -\frac{1}{2} + 2\times y\]

OpenStudy (anonymous):

then do 2 times 11/2 on the right hand side of the equation

OpenStudy (anonymous):

okay so the awnser is -1+2y=-11

OpenStudy (anonymous):

check your signs!

OpenStudy (anonymous):

oh its -1-2y=-11

OpenStudy (anonymous):

good, very good catch!

OpenStudy (anonymous):

okay thanks for your help